.

Neon gas is used in manufacturing to create signs. If you run an electric current through the gas,

it will glow in neon colors. Unfortunately, when you run electricity through the gas, its

temperature increases. If the temperature starts at 50⁰C, it will raise to about 150⁰C after the

sign is on for 4 hours. The gases original volume was 5 L. What is the volume of the gas after

the sign is on for 4 hours


1
Expert's answer
2022-03-28T11:02:31-0400

Solution

Let temperature is constat then final volume

"V'=\frac{VT'}{T}\\=\frac{5*150}{50}\\=15L"